OpenStudy (anonymous):

Under certain conditions argon (Ar) gas diffuses at a rate of 3.2 centimeters per second. Under the same conditions, an unknown gas diffuses at a rate of 4.5 centimeters per second. What is the approximate molar mass of the unknown gas? A.10 g/mol B. 20 g/mol C. 28 g/mol D. 56 g/mol

OpenStudy (aaronq):

Graham's law \[\frac{ rate 1 }{ rate 2 }=\sqrt{\frac{ Molarmass2 }{ Molarmass1 }}\] notice how the placement is inversed across the equal sign